two families attend the town ice cream social every summer. this year one family paid $9.75 for 3 adult tickets and 5 children tickets. Another family paid $10.75 for 5 adult tickets 1child ticket.Find the price of the child's tickets.
Answer by mananth(16949)

Adult ticket cost x
Child ticket cost------------ y
3x+5y=9.75.............1
5x+y=10.75 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 5
-3x-5y= -9.75
25x+5y=53.75
Add the two equations
22x=44
/22
x=2=$2.00 Adult ticket cost
plug value of x in (1)
3x+5y=9.75
6+5y=9.75
5y=9.75 -6
5y=3.75
y= 3/4=$0.75 Child ticket cost